OpenStudy (anonymous):

Mike has $1.15 in dimes and nickels. If he has 5 more nickels than dimes, how many of each coin does he have?

OpenStudy (alexandervonhumboldt2):

1.15$=115 cents. 1 dime=10 cents 1 nickel = 5 cents 10k+5n=115 10k+5(k+5)=115 solve this system

OpenStudy (alexandervonhumboldt2):

10k+5n=115 10k+5(k+5)=115 solve this system